# Machine Vision-Based Assessment of Fall Color Changes and its Relationship with Leaf Nitrogen Concentration

### v2

@@ Abstract
Apple trees, being deciduous, shed leaves each year. This process is preceded by the change in leaf color from green to yellow (also known as senescence) during the fall season. The rate and timing of color change are affected by factors including nitrogen (N) concentration in leaves. The green color of leaves and the speed at which it turns yellow during senescence are highly dependent on chlorophyll content, which in turn depends on nitrogen concentration in the leaves. The assessment of leaf color during this period can therefore provide important information on the nitrogen status of apple trees. This study focused on a machine vision-based system to quantify the timing and change in leaf color during the fall and correlate that information to leaf nitrogen content. The color and 3D image dataset used in this study was collected over five weeks in the fall of 2021 and 2023 at a commercial orchard (tall spindle architecture) using a ground-vehicle-based stereo vision sensor. First, the point cloud obtained from the sensor was used to segment the tree canopies using color and depth thresholding methods in the natural orchard environment. Then, the color information of the segmented canopy area was quantified using a custom-defined metric, yellowness index (a normalized ratio of yellow to green foliage in the tree) that varied from $-1$ to $+1$ ($-1$ being completely green and $+1$ being completely yellow), which gives the proportion of yellow leaves on a tree. A K-means-based method and a gradient boosting method were used to estimate the yellowness index. The gradient boosting method proposed in this study was found to be superior to the K-means-based method, achieving an $R^{2}$ of 0.72 in estimating the yellowness index. The results showed that the metric was able to capture the gradual color transition from green to yellow over the study period. It was also observed that the trees with lower nitrogen showed the color transition to yellow earlier than the trees with higher nitrogen. The onset of color transition during both years aligned with the 29th week post-full bloom. This critical timing could be used for conducting nitrogen status analysis on apple trees using machine vision, enabling more precise and timely assessment of nutrient levels and facilitating targeted fertilization strategies in orchard management.
@@
Leaf Nitrogen , Machine Vision in Agriculture , Point Cloud Segmentation , Precision Nitrogen Management , Machine Learning in Agriculture
@@
[inst3]organization=Mid-Columbia Agricultural Research and Extension Center,addressline=, Oregon State University, city=Hood River, postcode=97031, state=OR, country=USA
@@ 1 Introduction
Currently, tree fruit orchards are managed at at block or field level providing inputs such as water and nutrients uniformly to hundreds or even thousands of trees in the block without addressing the needs of individual trees. This coarse management of crops often leads to sub-optimal inputs thus negatively impacting tree growth, yield, and fruit quality (Klein et al., 1989; Neilsen et al., 2009). Individual tree-level management in tree fruit crops such as apple orchards presents a promising frontier for significantly enhancing fruit yield and quality addressing the unique needs of individual trees. However, managing these orchards (Figure 1) at the individual tree level is challenging, particularly in terms of assessing the needs of individual trees and making precise management decisions. Each tree within an orchard is unique with its specific needs for water, nutrients, and other inputs (Aggelopoulou et al., 2010; James A. Taylor et al., 2007). This variability can be caused by multiple factors including spatial variation of soil properties like texture, depth, slope, and microbial activities, and variation in microclimate within orchards (Aggelopoulou et al., 2010; Umali et al., 2012; Cho, 2010). Effective input management at the individual tree level could be possible if crucial factors such as nitrogen and water availability to individual trees can be accurately assessed.
@@
Nitrogen (N) is a major input to apple trees that impacts growth, health, and fruit quality (Sanchez et al., 1995) . Traditionally, the most accurate approach to assess tree N status involves laboratory-based leaf and soil nitrogen measurement, which provides a quantitative baseline for assessing the need for nitrogen application. However, these methods demand significant time and expense (Baret and Fourty, 1997) , limiting sampling density and the possibility of individual tree-level management in commercial orchards. As a result, orchard management often relies on data collected from sparingly located sample trees or sampling locations, lacking precision due to inherent spatial variations within orchards (Ferguson and Triggs, 1990; Wulfsohn et al., 2012; Arnó et al., 2017)
@@
These limitations of both laboratory-based and visual assessment methods underscore the need for a rapid, reliable, and more comprehensive approach to N status evaluation (Cheng and Schupp, 2004) at individual tree level. A machine vision-based system offers a promising solution by automating and standardizing the quantification of visual parameters across entire orchards. This approach could provide objective, quantifiable data that can be correlated with nitrogen status through various models. By enabling more precise, tree-level management in commercial orchards, such a system has the potential to overcome the sampling and subjectivity issues inherent in current methods while offering scalability for broader applications.
@@
Machine vision techniques have been extensively used to acquire various canopy-level features and stresses in fruit crops (Wang et al., 2023; Paudel et al., 2022; Wachs et al., 2010; Wang et al., 2018; Brown et al., 2024). . Specifically, many recent studies have reported machine vision-based tools for estimating canopy density(Mahmud et al., 2021; Paudel et al., 2023), tree trunk cross-sectional area (Wang et al., 2023), and canopy color (Naschitz et al., 2014) in fruit trees.
@@
Leaf color is one of the key canopy characteristics that is correlated to nitrogen content (Ali et al., 2012; Treder et al., 2016; Ye et al., 2020). This correlation is primarily due to the presence of chlorophyll in leaves. Chlorophyll, the pigment responsible for photosynthesis, contains nitrogen as a key component of its molecular structure. It reflects light primarily in the green spectrum, which is why leaves appear green to our eyes. In deciduous trees like apples, a significant color change occurs as the growing season progresses towards fall. This change is primarily due to the relocation of nitrogen from leaves to woody regions for storage and to support spring growth (Murneek and Logan, 1932; Spencer, 1973). As chlorophyll breaks down and moves back into the tree, the leaves transition from green to yellow.
This color transition provides a visual indicator of the tree’s nitrogen status. A greener canopy with vigorous growth is often correlated with a tree having high nitrogen, while a yellower and less vigorous canopy is often associated with trees with low nitrogen (Raese et al., 2007; Lee, 2002). Trees with very high nitrogen tend to maintain their green color longer into the season. The canopy color during this critical transition period can provide a unique and accurate indicator of tree nitrogen status.
While various RGB, multispectral, and hyperspectral imaging techniques have been used in the past to quantify the color to estimate nitrogen or chlorophyll content in leaves, most studies have focused on crops in controlled environments and at the individual leaf level. This focus limits their applicability to outdoor orchard conditions. Our study addresses this gap by using the images in natural orchard environments, and at the whole canopy level. We also aim to quantify canopy color over a senescence period when color difference in the tree foliage is more apparent, which can be used to assess nitrogen status of each tree in commercial orchards, potentially enabling precision management practices at the individual tree level.
Specifically, this study tracked apple tree canopies over five weeks in fall using RGB-D cameras, quantifying their color within an outdoor orchard environment and related the color at those critical periods of transition to the nitrogen levels in the corresponding trees. Specifically, the study focused on the following two research objectives:
@@
    To develop a machine vision-based technique to quantify foliage color during late fall in an outdoor apple orchard.
@@
    To use the foliage color to differentiate between trees with varying foliage nitrogen concentration.
The rest of the paper is structured into three sections. Section 2 details the data collection and processing methodologies. Section 3 delves into the results, offering insightful interpretation and pertinent discussion surrounding their emergence. Finally, Section 4 summarizes the key findings with discussions about the current limitations and future potentials.
@@ 2 Materials and Methods
The overall data collection and processing method used in this study is shown in Figure 2. Further detail about each step has been explained in the subsequent sections. The study consisted of collecting RGB-D images over multiple weeks in fall 2021 and 2023. The point cloud data of the tree canopies were then segmented into yellow and green foliage and a metric - yellowness index was defined and estimated to quantify the color of the trees. The yellowness indices of the tree during different weeks was then correlated to the leaf nitrogen content for each year.
Figure 2: Overall Data collection and processing method for the study. Each light colored box represent a distinct process and has been described further in subsequent sections.
@@ 2.1 Data Collection
The dataset (color and 3D images) for this study was collected over five weeks in Fall, 2021 (from October 13 to November 14) and Fall, 2023 (from October 20 to November 17). These images were collected in a Jazz commercial orchard (Yakima Valley Orchard) in Prosser, WA (Figure 3). Start day of the data collection was determined by when the leaves started to change the colors from green to yellow (Figure 4), also known as senescence. A total of 200 sample trees were scattered across 17 rows of the test orchard, with every 3rd tree in each row of trees taken as a sample. The data was collected using a commercial stereo-vision-based RGB-D camera (Zed2i, Stereolabs, Paris, France) mounted on a utility vehicle (Figure 3b). The trees were trained to a vertical structure using a tall spindle canopy architecture (Figure 1). The camera was mounted  7 feet above the ground to ensure the majority of the canopy was visible. An SVO file (Stereolabs file format) with all the metadata including the camera parameters, images, and sensor data was recorded for each row of trees. Images (pixel resolution of 2208 x 1242) and point cloud (maximum depth of 10m) for each tree were later extracted based on timestamp matching between frames and user input. During processing, the RGB and co-registered point cloud were identified for each test tree.
Figure 3: Study site and sensing setup; (a) A satellite view of the test plot. The red outline shows the location of the plot used in this study., and b) Ground vehicle with a camera mounted on top (zoomed portion shows camera and axes orientations),
Figure 4: Images of the same tree during different weeks of the data collection. The dates below represent the exact dates when the data was collected in 2023. The foliage can be seen gradually changing from green to yellow.
@@ 2.2.1 Foreground Tree Segmentation
The initial data acquired through the stereovision sensor was noisy and required further refinement. An initial step involved applying a color threshold of 153 to the blue channel to filter out pixels corresponding to the sky. Subsequently, a distance threshold of 3 meters along the Z-axis (depth axis), was implemented. The optimal threshold values were determined experimentally through trials and visual inspections. Consequently, the resulting point cloud consisted of points mostly from the foreground tree, with some residual points from the ground. To eliminate the undesirable ground points, the points within 50 centimeters above the lowest x-value (along the height of the tree) were excluded. The resulting point cloud was down-sampled by selecting every tenth point, effectively reducing its size while preserving essential information for subsequent analysis.
@@ 2.2.2 Color Segmentation and clustering
The point cloud obtained from step 2.2.1 consisted of points from the foreground tree, and could be broadly categorized into yellow leaves, green leaves, and the trunk. Data collection spanned multiple weeks, capturing the tree’s seasonal shift during the fall color change, depicted in Figure 4. consisted of points from the foreground tree, and could be broadly categorized into yellow leaves, green leaves, and the trunk. Data collection spanned multiple weeks, capturing the tree’s seasonal shift during the fall color change, depicted in Figure (Figure 6a) and CIE-Lab* (Figure 6b and c) color spaces were used to reduce the effect of varying intensity. In the HSV space, hue remains unaffected by varying light intensity, while in the CIE-Lab* space, the a* and b* values correspond respectively to red-green and blue-yellow colors. RGB images captured by the RGB-D sensor were converted to both the CIE-La*b* (Lab) space, using a standard D65 illuminant, and the HSV color space. These conversions were performed using the scikit-image library (Van der Walt et al., 2014).
The transition of the trees from green to yellow color was noticeable in both HSV and La*b* space (Figures 6a and 6b,c). A probability density plot for the hue space across the weeks illustrated a decline from approximately 110 degrees in week 1 to about 50 degrees in week 6, signifying the shift from green to yellow. A similar trend was visible in the probability density plots for a* and b* in the La*b* space. Both a* and b* showed a slight increase as the weeks progressed, corresponding to a color change from green to yellow.
@@
Output: Clustered Point cloud–3 groups, Green, Yellow, Trunk
1 $n$ = initial number of cluster (integer)
2 K-Means Clustering using $a$, $b$ values into $n$ clusters ($c$)
3 Green cluster($c_{g}$) = 0; Yellow cluster($c_{y}$) = 0; Trunk cluster ($c_{t}$) = 0
@@
5for $i$ in range($n$) do
6      if $cg_{amin}<c_{i}(a)<cg_{amax}$ and $cg_{bmin}c_{i}(b)<cg_{bmax}$ then
7           $c_{g}$ = $c_{g}$+$c_{i}$;
8           else if $cy_{amin}<c_{i}(a)<cy_{amax}$ and $cy_{bmin}c_{i}(b)<cy_{bmax}$ then
9                $c_{y}$ = $c_{y}$+$c_{i}$
10                else if $ct_{amin}<c_{i}(a)<ct_{amax}$ and $ct_{bmin}c_{i}(b)<ct_{bmax}$ then
11                    $c_{t}$ = $c_{t}$+$c_{i}$

### v3

# Machine Vision-Based Assessment of Fall Color Changes in Apple Leaves and its Relationship with Nitrogen Concentration

@@ Abstract
Apple(Malus domestica Borkh.) trees are deciduous, shedding leaves each year. This process is preceded by a gradual change in leaf color from green to yellow as chlorophyll is degraded prior to abscission. The initiation and rate of this color change are affected by many factors including leaf nitrogen (N) concentration. We predict that leaf color during this transition may be indicative of the nitrogen status of apple trees. This study assesses a machine vision-based system for quantifying the change in leaf color and its correlation with leaf nitrogen content. An image dataset was collected in color and 3D over five weeks in the fall of 2021 and 2023 at a commercial orchard using a ground vehicle-based stereovision sensor. Trees in the foreground were segmented from the point cloud using color and depth thresholding methods. Then, to estimate the proportion of yellow leaves per canopy, the color information of the segmented canopy area was quantified using a custom-defined metric, “yellowness index” (a normalized ratio of yellow to green foliage in the tree) that varied from -1 to +1 (-1 being completely green and +1 being completely yellow). Both K-means-based methods and gradient boosting methods were used to estimate the yellowness index. The gradient boosting based method proposed in this study was better than the K-means-based method (both in terms of computational time and accuracy), achieving an $R^{2}$ of 0.72 in estimating the yellowness index. The metric was able to capture the gradual color transition from green to yellow over the study duration. Trees with lower leaf nitrogen showed the color transition to yellow earlier than the trees with higher nitrogen. The onset of color transition during both years occurred during the $29^{th}$ week post-full bloom (October 22 in 2021 and Nov 10 in 2023). This critical timing could be used for conducting nitrogen status analysis on apple trees using machine vision, enabling more precise and timely assessment of nutrient levels and facilitating targeted fertilization strategies in orchard management.
@@
Leaf Color Change , Machine Vision , Point Cloud Segmentation , Precision Nitrogen Management , Machine Learning
@@
[…]
@@ 1 Introduction
Water and nutrient applications in commercial tree fruit orchards are managed at a block or field level, providing treatment relatively uniformly to hundreds or even thousands of trees without considering the variability among individual trees. This homogeneous management of nutrients in tree fruit crops often leads to sub- and/or super-optimal inputs, with negative impacts on tree growth, fruit yield, and fruit quality (Klein et al., 1989; Neilsen et al., 2009). Managing these inputs at the individual tree level (i.e. addressing the unique needs of individual trees) is a promising concept for improving tree vigor, fruit yield and quality. However, managing orchards (Figure 1) at the individual tree level is challenging, particularly in terms of assessing the needs of individual trees and making precise management decisions. Each tree within an orchard is unique with its specific needs for water, nutrients, and other inputs (Aggelopoulou et al., 2010; James A. Taylor et al., 2007). For eg., Aggelopoulou et al. (2010) reported spatial variability with a coefficient of variation (CV) of up to 82% for bloom and 40% for yield within the same block. The variability among trees can be caused by multiple factors including spatial variation of soil properties like texture, depth, slope, or microbial activities, variation in microclimates within orchards, and differences in tree training and pruning (Aggelopoulou et al., 2010, 2011; Umali et al., 2012; Cho, 2010; Parkin, 1993). Effective nutrient management at the individual tree level could be possible if these multiple factors can be accurately assessed for each individual tree.
@@
[…]
In addition to laboratory-based tests, experienced farmers and researchers often employ visual assessment of canopy features such as shoot length, canopy density, trunk diameter, overall tree growth, crop load, and leaf color to estimate N status.. While these observational methods leverage professional expertise, they suffer from subjectivity, and potential variability among assessors. The reliance on seasoned professionals further limits its widespread applicability in commercial orchards.
@@
These limitations of both laboratory-based methods and visual assessments by subject matter experts underscore the need for a rapid, reliable, and more comprehensive approach to N status evaluation (Cheng and Schupp, 2004) at the individual tree level. A promising alternative is a machine-vision system that automates and standardizes the quantification of visual parameters across entire orchards. This approach could provide objective, empirical data that may be correlated with tree nitrogen status through modeling methods. By enabling more precise, tree-level management in commercial orchards, such a system has the potential to overcome the sampling and subjectivity issues inherent in current methods while offering scalability for broader applications.
@@
Machine vision techniques have been extensively used to evaluate various canopy-level features and detect canopy stresses in fruit crops (Wang et al., 2023; Paudel et al., 2022; Wachs et al., 2010; Wang et al., 2018; Brown et al., 2024). More specifically, many recent studies have reported machine vision-based tools for estimating canopy density(Mahmud et al., 2021; Paudel et al., 2023), tree trunk cross-sectional area (Wang et al., 2023), and canopy color (Naschitz et al., 2014) in fruit trees.
@@
Leaf color is one of the key canopy characteristics that is correlated with nitrogen content(Ali et al., 2012; Treder et al., 2016; Ye et al., 2020). This correlation is due to the presence of four rings of N in the chlorophyll, the pigment in leaves responsible for light harvesting(Marschner, 2011). Chlorophyll reflects light primarily in the green spectrum, which is why plant tissue containing chlorophyll appears green. In deciduous trees like apples, a significant color change occurs from green to yellow as the growing season progresses towards fall, as the chlorophyll breaks down and mobile elements like N move back into the perennial structures of tree. This relocation of nitrogen from leaves to woody regions for storage will later support next seasons’ growth(Murneek and Logan, 1932; Spencer, 1973; Neilsen and Neilsen, 2002).A greener canopy with vigorous growth is often correlated with a tree having high nitrogen, while a yellower and less vigorous canopy is often associated with trees with low nitrogen (Raese et al., 2007; Lee, 2002). In addition, trees with high nitrogen tend to maintain their green color longer into the season providing a potential indication of tree nitrogen status during this critical transition.
Various RGB, multispectral, and hyperspectral imaging techniques have been evaluated to quantify the color and estimate nitrogen or chlorophyll content in apples (Campbell et al., 1990; Neilsen et al., 1995; Perry and Davenport, 2007). Most of these studies use devices like chlorophyllmeter (SPAD) or satellite images. While handheld devices like SPAD meters can be utilized in the field, they pose challenges for scaling assessments across entire orchards due to the need for data collection from individual leaves. On the other hand, satellite imagery offers broader insights but often lacks the resolution required to pinpoint individual trees, necessitating further computations for accurate matching. Our study overcomes these limitations by focusing on individual tree-level assessments, providing a scalable and efficient approach for evaluating apple tree health and productivity in orchards.
In this study, we investigated the relationship between N status and color degradation utilizing RGB-D camera during apple tree senescence, . and explored the potential to utilize this assessment for precision management practices at the individual tree level.
Specifically, the study had the following two objectives:
@@
    Develop a machine vision-based technique to quantify foliage color during late fall in an apple orchard.
@@
    Use the foliage color as an indicator to differentiate between trees with varying foliage leaf nitrogen concentration.
@@ 2 Materials and Methods
RGB-D images were collected over several weeks in fall of 2021 and 2023. These images were collected in a Jazz commercial orchard (Yakima Valley Orchard) near Prosser, WA, USA (Figure 3). The trees were trained to a vertical structure using a tall spindle canopy architecture (Figure 1). The point clouds of the tree canopies were then segmented into yellow and green foliage; a quantitative metric, termed the yellowness index, was created to estimate the color of the trees. The yellowness indices of the trees during different weeks were then correlated with the leaf nitrogen content for each year. The overall data collection and processing method used in this study is shown in Figure 2.
Figure 2: Overall data collection and processing steps used in the study; Each light-colored box represents a distinct process and has been described further in the subsequent sections.
@@ 2.1 Data Collection
The datasets (color and 3D images) for this study were collected over five weeks in fall, 2021 (from October 13 to November 14) and fall, 2023 (from October 20 to November 17), every week. The first image datasets were determined by when the leaves began to change color from green to yellow (Figure 4), The dataset was collected from a total of 200 trees across 17 rows of the test orchard, with every other 3rd tree in each row taken as a sample. The data were collected using a commercial stereovision-based RGB-D camera (Zed2i, Stereolabs, Paris, France) mounted on a utility vehicle (Figure 3b). The camera was mounted  2 meters above the ground to ensure that most of the canopy (Figure 4, part of tree above  0.3m from ground) was visible. An SVO file (Stereolabs file format) with all the metadata including the camera parameters, images, and sensor data was recorded for each row of trees. Images (pixel resolution of 2208 x 1242) and point clouds (maximum depth of 10m) for each tree were later extracted based on timestamp matching between frames and user input. During processing, the RGB and co-registered point cloud were identified for each test tree.
Figure 3: Study site and sensing setup; (a) An aerial view of the test plot. The red outline shows the location of the plot used in this study., and b) Ground vehicle with a camera mounted on top (zoomed portion shows camera and axes orientations)
Figure 4: Images of a sample apple tree acquired over the data collection period in 2023. The foliage can be seen gradually changing from green to yellow.
@@ 2.2.1 Foreground Tree Segmentation
The initial data acquired through the stereovision sensor were noisy and required further refinement. An initial step involved applying a color threshold of 153 to the blue channel to filter out pixels corresponding to the sky in the RGB color space. Subsequently, a distance threshold of 3 meters along the Z-axis (depth axis), was implemented. The optimal threshold values were determined experimentally through trials and visual inspections. Consequently, the resulting point cloud consisted of points mostly from the foreground tree, with some residual points from the ground. To eliminate the undesirable ground points, the points within 50 centimeters above the lowest x-value (along the height of the tree) were excluded. The resulting point cloud was down-sampled by selecting every tenth point, effectively reducing its size while preserving essential information for subsequent analysis.
@@ 2.2.2 Color Segmentation and Clustering
The point cloud obtained from step 2.2.1 consisted of points from the foreground tree, and could be broadly categorized into yellow leaves, green leaves, and the trunk. To quantify the color changes in trees’ over the duration of study, HSV (Figure 6a) and CIE-Lab* (Figure 6b and 6c) color spaces were used as they are less affected by the variation in intensities. In the HSV space, hue remains unaffected by varying light intensity, while in the CIE-Lab* space, the a* and b* values correspond respectively to red-green and blue-yellow colors. RGB images captured by the RGB-D sensor were converted to both the CIE-La*b* (Lab) space, using a standard D65 illuminant, and the HSV color space. These conversions were performed using the scikit-image library (Van der Walt et al., 2014).
The transition of the trees from green to yellow color was noticeable in both HSV and La*b* space (Figures 6a and 6b,c). A probability density plot for the hue space across the weeks illustrated a decline from approximately 110 degrees in week 1 to about 50 degrees in week 6 (hue expressed as degrees between 0 and 360, as shown in Figure 6a), signifying the shift from green to yellow. A similar trend was visible in the probability density plots for a* and b* in the La*b* space. Both a* and b* showed a slight increase as the weeks progressed, corresponding to a color change from green to yellow.
@@
Output: Clustered Point cloud–3 groups: Green ($c_{g}$), Yellow ($c_{y}$), Trunk ($c_{t}$)
1
2Define $n$ as the initial number of clusters (integer)
3 Perform K-Means clustering using $a*$, $b*$ values into $n$ clusters ($C=\{c_{1},c_{2},...,c_{n}\}$)
@@
5Initialize Green cluster ($c_{g}$), Yellow cluster ($c_{y}$), and Trunk cluster ($c_{t}$) as empty sets:
6 $c_{g}=\emptyset$; $c_{y}=\emptyset$; $c_{t}=\emptyset$
7
8Define $a*$ and $b*$ thresholds for Green cluster: $[cg_{amin},cg_{amax}]$, $[cg_{bmin},cg_{bmax}]$
9 Define $a*$ and $b*$ thresholds for Yellow cluster: $[cy_{amin},cy_{amax}]$, $[cy_{bmin},cy_{bmax}]$
10 Define $a*$ and $b*$ thresholds for Trunk cluster: $[ct_{amin},ct_{amax}]$, $[ct_{bmin},ct_{bmax}]$
11
12for $i\in\{1,2,...,n\}$ do
13      if $cg_{amin}\leq c_{i}(a)\leq cg_{amax}$ and $cg_{bmin}\leq c_{i}(b)\leq cg_{bmax}$ then
14           $c_{g}=c_{g}\cup c_{i}$
           // Merge cluster $c_{i}$ into Green cluster
15
16           else if $cy_{amin}\leq c_{i}(a)\leq cy_{amax}$ and $cy_{bmin}\leq c_{i}(b)\leq cy_{bmax}$ then
17                $c_{y}=c_{y}\cup c_{i}$
                // Merge cluster $c_{i}$ into Yellow cluster
18
19                else if $ct_{amin}\leq c_{i}(a)\leq ct_{amax}$ and $ct_{bmin}\leq c_{i}(b)\leq ct_{bmax}$ then
20                     $c_{t}=c_{t}\cup c_{i}$
                     // Merge cluster $c_{i}$ into Trunk cluster
21